OpenStudy (anonymous):

Factor 6x^4 - 4x^2 + 8x Factor 8p^3 + 12p^2 + 4p Factor 12k^3 + 6k^2 - 18k It would be helpful if you show how you get the answers, please and thank you very much.

myininaya (myininaya):

the gcf of 6x^4,4x^2,8x is 2x so we have 2x(3x^3-2x+4)

OpenStudy (anonymous):

im aware of all the GCFs but how do i factor them afterwards

myininaya (myininaya):

it is factored

myininaya (myininaya):

we just factored it i mean

OpenStudy (anonymous):

No... How do you factor the problem completely

OpenStudy (anonymous):

8p^3 + 12p^2 + 4p =4p(2p^2+3p+4) GCF the trinomial in this one and the one that m'naya did do not factor over the integers

OpenStudy (anonymous):

oh okay so that is the final answer

myininaya (myininaya):

you can try syntheic division to find a zero of 3x^3-2x+4

myininaya (myininaya):

but that only works if it has rational roots

OpenStudy (anonymous):

12k^3 + 6k^2 - 18k =6k(2k^2+k-3) GCF =6k(2k+3)(k-1) there is a method for factoring the trinomial, but since 2 and 3 are prime trial and error is what i used; i'll show you if you want

OpenStudy (anonymous):

please do..

OpenStudy (anonymous):

ac=2*(-3)=-6 then factors of ac that sum to b are -2*3=-6 and -2+3=+1 so we can use -2 and 3 as coefficients to rewrite the +k term =2k^2 -2k +3k -3 then we factor by grouping (take GCF from first two and last two terms) =2k(k-1)+3(k-1) to finish factoring by grouping we factor out the k-1 as a GCF =(k-1)(2k+3)

OpenStudy (anonymous):

a, b, and c refer to \[ax^2+bx+c\]

OpenStudy (anonymous):

You know that factoring and multiplication are inverse operations. the method takes you to the step following a FOIL: (k-1)(2k+3)=2k^2+3k-2k-3 compare the above line to my previous post; they are the same except that the middle two terms are commuted.